# Solving Inverse Problems in Stochastic Self-Organising Systems through Invariant Representations

### v1

@@ 4 Experiments and Results
In this section, we present the results of applying the method to two self-organising systems: a reaction-diffusion system known as the Gray-Scott model [45, 46], and Schelling’s agent-based model of segregation [1]. We chose these two distinct systems, both in terms of scientific domain (biophysical vs. social) and the underlying mathematical model (PDE vs. ABM), to highlight the generality of the method. We first introduce the models, then formulate them as inverse problems, and subsequently present the results of applying the method. Finally, we evaluate the method on two non-simulated patterns: an ocellated lizard and a zebra skin pattern.
Hardware details. Experiments have been run on a single machine with a AMD Ryzen Threadripper 3990X CPU and a NVIDIA GeForce RTX 4090 GPU. Reproducibility. The experiments can reproduced using the scripts found here: https://github.com/enajx/InverseStochastic/experiments.
4.1 Reaction-Diffusion system
Gray-Scott is a reaction-diffusion model used to study pattern formation in developmental systems [45, 46]. It describes how spatially distributed concentrations of two interacting substances– or morphogens–evolve over time. Morphogens play a critical role in developmental biology by guiding cell differentiation and tissue patterning through their concentration gradients. $u$ represents a nutrient or precursor substance, while $v$ acts as an autocatalyst that promotes its own production while consuming $u$. The interaction between $u$ and $v$, combined with their ability to diffuse through space, leads to the emergence of complex patterns such as spots, stripes, or labyrinths. Through this interplay, the Gray-Scott model provides insights into how local reactions and diffusion contribute to the self-organisation of biological structures. It is formalised as two partial differential equations:
$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+f(1-u)$ $\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(f+k)v$ $u$ and $v$ are the chemical concentrations. - $D_{u}$ and $D_{v}$ are the diffusion rates for $u$ and $v$. - $f$ is the feed rate of $u$. - $k$ is the kill rate of $v$. - $\nabla^{2}$ is the Laplacian representing diffusion.
@@ 4.1.1 Results Reaction-Diffusion
In order to find the unknown parameters for the reaction-diffusion system, we follow the algorithm described in Algorithm 1. At each optimisation run, we first generate a target pattern for a set of parameters $\{D_{u},D_{v},k,f\}_{target}\in\mathbb{R}^{4}$ by simulating Gray-Scott partial differential equations 4.1 from a random initial state $s_{0}\sim\mathcal{S}$ for 1000 steps. The resulting pattern is embedded onto $z_{target}$ with the vision model CLIP. $z_{target}$ serves as the target embedding for the loss throughout the optimisation process. Then, the optimiser samples a population of candidate solutions, $\{D_{u},D_{v},k,f\}_{i}\in\mathbb{R}^{4}$, used to simulate the dynamics of reaction-diffusion. The generated patterns $y_{i}$ are then passed through the embedding model, resulting in $z_{i}$. We compare each $z_{i}$ against the target embedding $z_{target}$ using cosine similarity $cos(z_{i},z_{target})$. The optimisation is iteratively run until it converges to a solution that maximises the cosine similarity with respect to the target embedding $z_{target}$. Results for three sets of reaction-diffusion parameters and its corresponding training curves are shown in Figure 5.
@@
Figure 5: Results for Gray-Scott reaction-diffusion system. (a): A selection of three target patterns $y_{i}$ (left column), and three reconstructions $y_{i}^{\prime},y_{i}^{\prime\prime},y_{i}^{\prime\prime\prime}$ of the causal parameters found by the inverse method on three independent training runs. The method reconstructs patterns with features that qualitatively match the targets. (b): Loss curves showing the average loss and variance of 10 training runs for 12 different patterns. (c): Histogram showing the mean-square errors between the causal $\theta$ and predicted parameters $\theta^{\prime}$; error bars indicate one standard deviation of the 10 training runs averaged for all the 12 target patterns.
@@
Figure 10 in the Appendix shows the results of 10 independent training runs on a set of 12 target patterns. The loss curves and histogram in Figure 5 reflect results for the full set of twelve target patterns. Figure 12 shows the cosine distance matrix between CLIP embeddings of the set of target patterns used.
4.2 Schelling’s Model of Segregation
@@
The second system we use to demonstrate the method is Schelling’s segregation model [1], a classic computational model used to study the emergence of spatial patterns in social systems. It illustrates how individual preferences for neighbourhood composition can lead to large-scale segregation, even with mild preferences. This model has been highly influential in computational social science, specifically contributing to the understanding of residential segregation and its implications for urban planning and social dynamics. Schelling’s model has a historical significance in computational social science, as it was one of the first to demonstrate how simple local rules can produce complex emergent behaviours, showing the potential of agent-based models to investigate social systems [16, 47]. It is for this latter reason that we chose to demonstrate our method with it.
@@
Our implementation consists of two classes of agents, distributed on a spatial grid. Each agent evaluates its satisfaction based on the proportion of similar neighbours in its local neighbourhood. Unsatisfied agents with their current location will asynchronously relocate to a random location to improve their satisfaction. Schelling’s model leads to the emergence of complex segregation patterns that match those observed in urban environments.
@@
$S_{i}=\begin{cases}1&\text{if }\frac{N_{\text{similar}}}{N_{\text{total}}}\geq T\\ 0&\text{otherwise}\end{cases}$ $S_{i}$ is the satisfaction of agent $i$. - $N_{\text{similar}}$ is the number of similar neighbours. - $N_{\text{total}}$ is the total number of neighbours. - $T$ is the agent’s tolerance threshold, $T\in[0,1]$.
@@ 4.2.1 Results - Schelling Segregation Model
In this experiment, the same optimisation methodology described in 1 is applied to find the tolerance $T$ from different Schelling segregation patterns. We use a grid size of 100 by 100, and an occupation density of 90%. The results of the experiment are presented in Figure 6.
Figure 6: Results for Schelling segregation model. (a): Schelling runs with tolerance parameter $\mathbf{T}\in\{0.1\mathbf{i}\mid\mathbf{i}\in\{1,2,\dots,8\}\}$. For small values of $T$ all agents converges with minimal or no migration since agents are satisfied with diverse neighbours, in middle range of $T$ (3-6) segregation patterns become more pronounced, while for large values of $T$ the system doesn’t converge since there are not enough empty cells to allocate all agents under the defined satisfaction threshold. (b): Training curve showing the average loss and variance for target patterns associated to each system tolerance. (c): Histogram showing the mean-square errors between the causal $\theta$ and predicted parameters $\theta^{\prime}$; error bars indicate one standard deviation of the 10 training runs for each tolerance target.
The method successfully finds the inverse parameters. The curves and histogram of Figure 6 indicate the mean and standard deviation of 10 independent training runs. The patterns resulting from each training run are shown in the Appendix, Figure 11. The cosine distance matrix of the CLIP embeddings of the Schelling patterns is shown in the Appendix Figure 13.
@@ 4.3 Natural patterns
In this section, we evaluate the method on natural patterns–rather than simulated ones, as in the previous experiments. First, on the skin patterning from an ocellated lizard. We provide the method with a target image obtained from Manukyan et al. [8] and search for the unknown parameters of a Gray-Scott reaction-diffusion model to reconstruct the patterns following the same methodology described in 1. The target image and the reconstructed one are shown in Figure 7.
@@
Identifying unknown parameters underlying natural patterns is a significantly more challenging problem as there is no guarantee that the chosen model can generate the target patterns. We often lack an exact model of the underlying process. In such cases, one can approach the problem by including the functional form of the model itself and have the optimiser predict the model’s functional form and parameters–an optimisation technique known as symbolic regression [48]. However, the space of possible models is extremely large, and furthermore, we often have an intuition about the model. In such cases, we can conjecture a minimal model form that might be driving the process that generates the target pattern. We can then apply the inverse method to tune the free parameters of the model to reconstruct the natural target. If the model’s expressivity is insufficient to capture the pattern, we can iteratively complexify the model. This approach allows us to preserve a mechanistic understanding of the underlying process until the model captures the target pattern.
@@
$\frac{\partial u}{\partial t}$ $=D_{u_{x}}\frac{\partial^{2}u}{\partial x^{2}}+D_{u_{y}}\frac{\partial^{2}u}{\partial y^{2}}-uv^{2}+f(1-u)$ $\frac{\partial v}{\partial t}$ $=D_{v_{x}}\frac{\partial^{2}v}{\partial x^{2}}+D_{v_{y}}\frac{\partial^{2}v}{\partial y^{2}}+uv^{2}-(f+k)v$ $D_{u_{x}}$ and $D_{u_{y}}$ are the diffusion rates for $u$ in the $x$- and $y$-directions. - $D_{v_{x}}$ and $D_{v_{y}}$ are the diffusion rates for $v$ in the $x$- and $y$-directions.
Figure 8: Reconstruction of a zebra pattern. Left: Photograph of an zebra skin pattern [49]. Right: Pattern generated by the anisotropic reaction-diffusion model for the parameters found by the method. Cosine similarity between target and reconstruction $0.842\pm 0.018$, for 100 evaluations.
All other terms remain the same as in the standard Gray-Scott model. This anisotropic modification allows for spatial variation in the diffusion dynamics, introducing a directional bias that can generate the stripe patterns required for the zebra pattern. As a result, the anisotropic Gray-Scott model successfully produces the desired pattern, as shown in Figure 8.
Along with the other approaches discussed in Section 2, the expressivity bottleneck of the model exemplifies how our method can guide the search for mechanistic models. While the method can simply fail due to the optimisation being challenging, by iteratively expanding a conjectured model until it achieves an expressivity level compatible with observed data, we can use the method to potentially identify missing mechanisms and refine our understanding of pattern formation in nature.
@@ 5 Conclusion
In this work, we have presented a method for solving inverse problems in stochastic self-organising systems by leveraging invariant visual representations from embedding models–without the need to handcraft metric functions to capture the visual similarities between patterns. The key idea is to shift the optimisation from pixel space to an embedding space where perceptually similar patterns are mapped to nearby points, allowing the recovery of unknown parameters, even those resulting in stochastic patterns. Unlike existing techniques in inverse modelling, the method addresses the issue of stochasticity in the observable space, rather than in the causal space. We have demonstrated the method on two distinct self-organising models: a reaction-diffusion system and an agent-based segregation model, both exhibiting stochasticity in the observation space. For the reaction-diffusion model, we have shown that the approach can be applied to real biological patterns: an ocellated lizard and a zebra skin pattern. Finally, we have illustrated how the framework can guide model testing: when a conjectured model fails to generate the target, it suggests that the model’s expressivity is insufficient and extra terms are needed, a practical path for iterative refinement in mechanistic modelling. In summary, the method provides a simple yet effective technique for theorists and experimentalists to investigate and control self-organising stochastic systems.

### v2

@@ 4 Experiments and Results
In this section, we present the results of applying the method to three self-organizing systems: a reaction-diffusion system known as the Gray-Scott model [38, 26], a model of embryonic development [8], and Schelling’s agent-based model of segregation [43]. We chose these three distinct systems, both in terms of scientific domain (physics, biology and sociology) and the underlying mathematical model (PDE vs. ABM), to highlight the generality of the method. We first introduce the models, then formulate them as inverse problems, and subsequently present the results of applying the method. Finally, we evaluate the method on two non-simulated patterns: an ocellated lizard and a zebra skin pattern.
Hardware details. Experiments have been run on a single machine with a AMD Ryzen Threadripper 3990X CPU and a NVIDIA GeForce RTX 4090 GPU. Reproducibility. The codebase including the algorithm implementation and scripts used to run the reported experiments can be found in https://github.com/enajx/InverseStochastic.
4.1 Physical system: Reaction-Diffusion system
Gray-Scott is a reaction-diffusion model used to study pattern formation in developmental systems [38, 26]. It describes how spatially distributed concentrations of two interacting substances–or morphogens–evolve over time. Morphogens play a critical role in developmental biology by guiding cell differentiation and tissue patterning through their concentration gradients. $u$ represents a nutrient or precursor substance, while $v$ acts as an autocatalyst that promotes its own production while consuming $u$. The interaction between $u$ and $v$, combined with their ability to diffuse through space, leads to the emergence of complex patterns such as spots, stripes, or labyrinths. Through this interplay, the Gray-Scott model provides insights into how local reactions and diffusion contribute to the self-organization of biological structures. It is formalised as two partial differential equations:
$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+f(1-u)$ $\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(f+k)v$ - $u$ and $v$ are the chemical concentrations. - $D_{u}$ and $D_{v}$ are the diffusion rates for $u$ and $v$. - $f$ is the feed rate of $u$. - $k$ is the kill rate of $v$. - $\nabla^{2}$ is the Laplacian representing diffusion.
@@ 4.1.1 Results Reaction-Diffusion
In order to find the unknown parameters for the reaction-diffusion system, we follow the algorithm described in Algorithm 1. At each optimization run, we first generate a target pattern for a set of parameters $\{D_{u},D_{v},k,f\}_{target}\in\mathbb{R}^{4}$ by simulating Gray-Scott partial differential equations 4.1 from a random initial state $s_{0}\sim\mathcal{S}$ for 1000 steps. The resulting pattern is embedded onto $z_{target}$ with the vision model CLIP. $z_{target}$ serves as the target embedding for the loss throughout the optimization process. Then, the optimizer samples a population of candidate solutions, $\{D_{u},D_{v},k,f\}_{i}\in\mathbb{R}^{4}$, used to simulate the dynamics of reaction-diffusion. The generated patterns $y_{i}$ are then passed through the embedding model, resulting in $z_{i}$. We compare each $z_{i}$ against the target embedding $z_{target}$ using cosine similarity $cos(z_{i},z_{target})$. The optimization is iteratively run until it converges to a solution that maximizes the cosine similarity with respect to the target embedding $z_{target}$. Results for three sets of reaction-diffusion parameters and their corresponding training curves are shown in Fig. 5.
@@
Figure 5: Results for Gray-Scott reaction-diffusion system. (a): A selection of three target patterns $y_{i}$ (left column), and three reconstructions $y_{i}^{\prime},y_{i}^{\prime\prime},y_{i}^{\prime\prime\prime}$ of the causal parameters found by the inverse method on three independent training runs. The method reconstructs patterns with features that qualitatively match the targets. (b): Loss curves showing the average loss and variance of 10 training runs for 12 different patterns. (c): Histogram showing the mean-square errors between the causal $\theta$ and optimized parameters $\theta^{\prime}$; error bars indicate one standard deviation of the 10 training runs averaged for all the 12 target patterns.
@@
Results for 10 independent training runs on a set of 12 target patterns are shown in Fig. 11. The loss curves and histogram in Fig. 5 reflect results for the full set of twelve target patterns. Fig. 14 shows the cosine distance matrix between CLIP embeddings of the set of target patterns used.
4.2 Biological system: Embryonic development
A blastocyst is the early-stage of the mammalian embryo. It is composed of three main components, an outer cellular layer, the trophoblast, which later in development becomes the placenta, and an inner cellular cluster, the embryoblast that develops into the organism itself, and a layer, the hypoblast, separating the embryoblast from the internal cavity; the cellular arrangement of the embryo at the blastocyst stage is shown in Fig. 6a. Due to the stochastic nature of morphogenesis, the exact position where the stem cells differentiate the embryoblast varies with each developmental instance.
The blastocyst simulation is based on Cang et al. [8] Cellular Potts Model (CPM), where the gene regulatory network model defined as a set of stochastic differential equations guide the cells’ differentiation. Unlike in the reaction-diffusion experiments, the blastocyst has a single stochastic target pattern, representing the observed embryogenic developmental process in mammalian organisms [4, 8]. The simulation is performed using Morpheus library and based on Morpheus [31]’s implementation.
4.2.1 Results - Embryonic development
In accordance with the embryonic development literature [8], a successful developmental process is one that leads to a cellular arrangement where the inner cell mass, or embryoblast, adheres to the outer layer, or trophoblast, while having a thin endodermic layer, or hypoblast, facing towards the inner cavity. An instance of this developmental pattern is shown in Fig. 6a. The inverse problem consists in finding a set of four Hill coefficients that modulate the expression of several transcription factors and proteins, acting as a complex genetic switch that guides development. Next to the target pattern, Fig. 6a shows the outcome of three independent optimization runs that correctly recover the parameters resulting in successful developmental patterns. For ten independent optimization runs, using CLIP embedding model to guide the optimization results in seven successful pattern recoveries–by contrast, using MSE as the optimization metric, results in only one correct developmental patterning. Results of embedding model and baseline metrics are shown in Table 1, and the obtained patterns for ten optimization runs are shown in Fig. 12.
Figure 6: Results for blastocyst model. (a): Target pattern and found solutions from three independent runs to recover the blastocyst parameters. (b): Training curve showing the average loss and variance for the target pattern over ten optimization runs. (c): Histogram showing the mean-square errors between the causal $\theta$ and optimized parameters $\theta^{\prime}$; error bars indicate one standard deviation of the 10 training runs for each tolerance target.
4.3 Social system: Schelling’s Model of Segregation
@@
The third system we use to demonstrate the method is Schelling’s segregation model [43], a classic computational model used to study spatial pattern emergence in social systems. It illustrates how individual preferences for neighbourhood composition can lead to large-scale segregation. Schelling’s model has been highly influential in computational social science, contributing to the understanding of segregation and its implications for urban planning and social dynamics. Schelling’s model has historical significance in computational social science, as it was one of the first to show how simple local rules can produce complex emergent behaviours, highlighting the potential of agent-based models for social systems [37, 9]. For this latter reason, we chose it to demonstrate our method.
@@
Our implementation consists of two classes of agents, randomly distributed on a spatial grid. Each agent evaluates its satisfaction based on the proportion of similar neighbours in its local neighbourhood. Unsatisfied agents with their current location will asynchronously relocate to a random location to improve their satisfaction. Schelling’s model leads to the emergence of complex segregation patterns that match those observed in urban environments.
@@
$S_{i}=\begin{cases}1&\text{if }\frac{N_{\text{similar}}}{N_{\text{total}}}\geq T\\ 0&\text{otherwise}\end{cases}$ - $S_{i}$ is the satisfaction of agent $i$. - $N_{\text{similar}}$ is the number of similar neighbours. - $N_{\text{total}}$ is the total number of neighbours. - $T$ is the agent’s tolerance threshold, $T\in[0,1]$.
@@ 4.3.1 Results - Schelling Segregation Model
In this experiment, the same optimization methodology described in Algorithm 1 is applied to find the tolerance $T$ from different Schelling segregation patterns. We use a grid size of 100 by 100, and an occupation density of 90%. The results of the experiment are presented in Figure 7.
Figure 7: Results for Schelling segregation model. (a): Schelling runs with tolerance parameter $\mathbf{T}\in\{0.1\mathbf{i}\mid\mathbf{i}\in\{1,2,\dots,8\}\}$. For small values of $T$, agents converge with minimal migration as tolerance threshold is quickly satisfied, in middle range of $T$ (3-6) segregation patterns become more pronounced, while for large values of $T$ the system does not converge as threshold is unattainable. (b): Training curve showing the average loss and variance for target patterns associated with each system’s tolerance. (c): Histogram of mean-square errors between the causal $\theta$ and optimized parameters $\theta^{\prime}$.
The method successfully finds the inverse parameters. The curves and histogram of Fig. 7 indicate the mean and standard deviation of 10 independent training runs. The patterns resulting from each training run are shown in the appendix Fig. 13.
4.4 Baselines comparison
Table 1: Loss values (median ± MAD across ten runs) for ten runs per target pattern under two embedding metrics (CLIP, Nomic), Fréchet Wavelet and the domain-specific metrics. Results ranked by z-score across metrics.
Reaction-Diffusion Blastocyst Schelling z-score
CLIP $0.038\pm 0.021$ $321\pm 0.097 0.0$$044\pm 0.033$ 1.08-1.08
Fréchet Wavelet $0.050\pm 0.032$ $333\pm 0.087 0.0$$040\pm 0.031$ 0.99-0.99
Nomic $0.042\pm 0.022$ $408\pm 0.028 0.0$$054\pm 0.041$ 0.75-0.75
Dissimilarity index $0.222\pm 0.060$ $483\pm 0.107 0.0$$040\pm 0.028$ $0.39$
Edge density $0.111\pm 0.045$ $0.517\pm 0.198$ $0.265\pm 0.226$ $0.61$
MSE $0.158\pm 0.082$ $0.507\pm 0.099$ $0.200\pm 0.129$ $0.63$
Orientation variance $0.126\pm 0.072$ $0.639\pm 0.119$ $0.299\pm 0.239$ $1.19$
We compare the performance of the embedding representations against six other baselines: mean squared error (MSE) for pixel-level comparison, Nomic, a second embedding model [36]; Fréchet Wavelet, a metric based on wavelet packet decomposition [50]; and a set domain-specific metrics: dissimilarity index, which measures spatial clustering patterns in agent-based models; edge density, quantifying boundary structures; and orientation variance, capturing directional pattern features. Results show that CLIP achieves the lowest z-score (-1.08), demonstrating the best overall performance across systems. While dissimilarity index performs optimally for its specific domain (i.e., segregation cluster), it performs poorly when evaluated in other patterns (z-score: 0.39). Other domain-specific metrics such as edge density and orientation variance show substantially worse overall performance (z-scores: 0.61 and 1.19 respectively). MSE ranks second to last (z-score: 0.63)–expectedly, pixel-to-pixel distance does not provide a good error signal to guide the optimization. The experiments results are shown in Table 1; variability is reported using median absolute deviation (MAD) given non-normal error distributions. Results demonstrate the effectiveness of invariant embeddings for handling stochasticity in the observable space without requiring domain-specific loss metrics.
@@ 4.5 Natural patterns
In this section, we evaluate the method on natural patterns–rather than simulated ones, as in the previous experiments. First, on the skin patterning from an ocellated lizard. We provide the method with a target image obtained from Manukyan et al. [29] and search for the unknown parameters of a Gray-Scott reaction-diffusion model to reconstruct the patterns following the same methodology described in 1. The target image and the reconstructed one are shown in Fig. 8.
@@
Identifying unknown parameters underlying natural patterns is a significantly more challenging problem as there is no guarantee that the chosen model can generate the target patterns. We often lack an exact model of the underlying process. In such cases, one can approach the problem by including the functional form of the model itself and have the optimizer predict the model’s functional form and parameters–an optimization technique known as symbolic regression [10]. However, the space of possible models is extremely large, and furthermore, we often have an intuition about the model. In such cases, we can conjecture a minimal model form that might be driving the process that generates the target pattern. We can then apply the inverse method to tune the free parameters of the model to reconstruct the natural target. If the model’s expressivity is insufficient to capture the pattern, we can iteratively complexify the model. This approach allows us to preserve a mechanistic understanding of the underlying process until the model captures the target pattern.
@@
$\frac{\partial u}{\partial t}$ $=D_{u_{x}}\frac{\partial^{2}u}{\partial x^{2}}+D_{u_{y}}\frac{\partial^{2}u}{\partial y^{2}}-uv^{2}+f(1-u)$ $\frac{\partial v}{\partial t}$ $=D_{v_{x}}\frac{\partial^{2}v}{\partial x^{2}}+D_{v_{y}}\frac{\partial^{2}v}{\partial y^{2}}+uv^{2}-(f+k)v$ - $D_{u_{x}}$ and $D_{u_{y}}$ are the diffusion rates for $u$ in the $x$- and $y$-directions. - $D_{v_{x}}$ and $D_{v_{y}}$ are the diffusion rates for $v$ in the $x$- and $y$-directions.
Figure 9: Reconstruction of a zebra pattern. Left: Photograph of a zebra skin pattern [28]. Right: Pattern generated by the anisotropic reaction-diffusion model for the parameters found by the method. Cosine similarity between target and reconstruction $0.842\pm 0.018$, for 100 evaluations.
All other terms remain the same as in the standard Gray-Scott model. This anisotropic modification allows for spatial variation in the diffusion dynamics, introducing a directional bias that can generate the stripe patterns required for the zebra pattern. As a result, the anisotropic Gray-Scott model successfully produces the desired pattern, as shown in Fig. 9.
Along with the other approaches discussed in Section 2, the expressivity bottleneck of the model exemplifies how our method can guide the search for mechanistic models. While the method can simply fail due to the optimization being challenging, by iteratively expanding a conjectured model until it achieves an expressivity level compatible with observed data, we can use the method to potentially identify missing mechanisms and refine our understanding of pattern formation in nature.
@@ 5 Conclusion
In this work, we have presented a method for solving inverse problems in stochastic self-organizing systems by leveraging invariant visual representations from embedding models–without the need to handcraft metric functions to capture the visual similarities between patterns. The key idea is to shift the optimization from pixel space to an embedding space where perceptually similar patterns are mapped to nearby points, allowing the recovery of unknown parameters, even those resulting in stochastic patterns. Unlike existing techniques in inverse modeling, the method addresses the issue of stochasticity in the observable space, rather than in the causal space.
We have demonstrated the method on three distinct self-organizing systems spanning physics, biology and computational sociology: a reaction-diffusion system, a developmental model of embryogenesis, and an agent-based segregation model, all exhibiting stochasticity in the observation space. For the reaction-diffusion model, we have shown that the approach can be applied to real biological patterns: an ocellated lizard and a zebra skin pattern. Finally, we have illustrated how the framework can guide model testing: when a conjectured model fails to generate the target, it suggests that the model’s expressivity is insufficient and extra terms are needed, a practical path for iterative refinement in mechanistic modeling. In summary, the method provides a simple yet effective technique for theorists and experimentalists to investigate and control self-organizing stochastic systems.